# Fast Estimation of the Composite Link Model for Multidimensional Grouped Counts

### v1

@@ Abstract
This paper presents a significant advancement in the estimation of the Composite Link Model within a penalized likelihood framework, specifically designed to address indirect observations of grouped count data. While the model is effective in these contexts, its application becomes computationally challenging in large, high-dimensional settings. To overcome this, we propose a reformulated iterative estimation procedure that leverages Generalized Linear Array Models, enabling the disaggregation and smooth estimation of latent distributions in multidimensional data. Through applications to high-dimensional mortality datasets, we demonstrate the model’s capability to capture fine-grained patterns while comparing its computational performance to the conventional algorithm. The proposed methodology offers notable improvements in computational speed, storage efficiency, and practical applicability, making it suitable for a wide range of fields where high-dimensional data are provided in grouped formats.
[…]
@@ 1 Introduction
The Composite Link Model (CLM) is an advanced framework designed in situations when it is necessary to link each observation with a linear function of more than one predicted value. This is the case when the aim is to model indirect observations of counts, and this particular setting will be the focus of this paper. Initially proposed by Thompson and Baker (1981), the CLM extends the capabilities of the Generalized Linear Model (GLM, McCullagh and Nelder, 1989; Nelder and Wedderburn, 1972). The CLM enables the incorporation of complex relationships and dependencies in count data, making it particularly useful in scenarios where direct observations are not feasible. This extension maintains the flexibility and interpretability of the GLM while addressing the unique challenges posed by indirect count data.
@@
In areas such as demography or epidemiology, this situation often arises when working with mortality tables, where death counts are aggregated over ages and/or years. Using this context as an illustration, the aim of this paper is to estimate the underlying latent mortality patterns across both ages and years, and potentially months and weeks. In the absence of prior knowledge, it is natural to exploit the inherent order in the data and impose smoothness to obtain unique estimates. This can be achieved by incorporating a roughness penalty into the associated likelihood function as described by Eilers (2007). Hence, throughout this paper, we refer to it as the Penalized Composite Link Model (PCLM). The resulting scoring algorithm is a modified version of the classic iterative reweighted least-squares (IRLS) for GLMs, incorporating an additional penalty and a revised design matrix to account for the grouping structure.
While this approach has been successfully applied in similar contexts (Rizzi et al., 2015, 2018), including with Bayesian estimation procedures (Lambert and Eilers, 2009; Lambert, 2011), the problem size in two- or higher-dimensional settings often becomes quite large, rendering the computational cost of maximizing the penalized likelihood function infeasible. In this paper we propose an alternative estimation procedure based on a reformulation of IRLS that resembles the EM-algorithm (Dempster et al., 1977). We alternate between computation of the latent distribution of counts from the current parameter estimates and estimating the parameters from the actual and estimated observations. This approach could be extremely useful in common situations where grouping structure in one dimension is constant over the other dimension (this is a common case in mortality tables where we observed aggregated death counts in coarse age groups and the aggregation structure is similar for all years and other dimensions).
Although it is well established that the EM algorithm can suffer from slow convergence (Ng et al., 2012), the computational and storage advantages of our approach stem from reformulating the whole algorithm as a Generalized Linear Array Model (GLAM, Currie et al., 2006). This reformulation not only mitigates the traditional slowness of EM but also efficiently handles large systems of equations, significantly improving both efficiency and scalability.
@@
The structure of this paper is as follows: following this Introduction, Section 2 provides an overview of the Penalized Composite Link Model, focusing on the new estimation methodology within the Generalized Linear Array Models framework, as well as methods for computing variance and standard errors. Section 3 illustrates the practical utility of the proposed approach through the analysis of two mortality datasets in two-dimensional and three-dimensional settings. Finally, Section 4 concludes by summarizing the main findings.
@@ 2 The Penalized Composite Link Model over multidimensional arrays
Suppose that we observe an array of counts $\bm{Y}=(y_{i,j})$ along with explanatory variables $\bm{x}_{1}=(1,\ldots,n_{1})$ and $\bm{x}_{2}=(1,\ldots,n_{2})$. We assume that $Y_{i,j}$, the random variable corresponding to the observed counts $y_{i,j}$, follows the Poisson distribution with mean $\mu_{i,j}$, i.e.  $Y_{i,j}\sim P(\mu_{i,j})$. Let $\bm{y}=\verb"vec"(\bm{Y})$ denote the vectorized form of the observed counts, and $\bm{\mu}$ represent the corresponding vector of expected values. In our setting, observed counts are the result of the contribution of several latent observations which we aim to estimate over a finer resolution of the covariates $\bm{x}_{1}$ and $\bm{x}_{2}$. In other words, while we observe $\bm{y}$, a vector of length $n_{1}n_{2}$, we aim to estimate $\bm{\gamma}=(\gamma_{11},\gamma_{12},\ldots,\gamma_{m_{1}m_{2}})$, $n_{d}<m_{d}$ $d=1,2$. The latent distribution is related to $\bm{\mu}$ via the composition matrix $\bm{C}$, which describes how the elements of the latent vector $\bm{\gamma}$ are combined to yield $\bm{\mu}$. The resulting Composite Link Model for Poisson aggregated counts can thus be written as follows:
@@
For our illustrative example, we will utilize demographic data on mortality. Here, the counts represent observed deaths, with $\bm{x}_{1}$ corresponding to ages and $\bm{x}_{2}$ to calendar years. In many cases, especially when dealing with smaller geographic areas or specific causes of death, the available data are often aggregated over broader age groups or multiple years. Our objective is to estimate mortality at a more granular level, specifically for individual ages and single calendar years, to provide finer insights into mortality patterns. In general, it is common for the aggregation pattern in each dimension of the array to be consistent across all dimensions. For instance, if death counts are aggregated into 5-year age intervals, this grouping pattern is typically applied systematically across all (grouped) years of death. Furthermore, when analyzing death counts, it is essential to recognize that the expected values in the Poisson distribution are the product of the underlying force of mortality and the population at risk, commonly referred as exposures: $\bm{y}\sim\mathcal{P}(\bm{e}\odot\bm{\mu})$. Exposures can be simply multiplied the underlying latent force of mortality over a finer resolution since, they are typically provided for single years of age and time, i.e. $\bm{e}$ represents the vectorized form of the $m_{1}\times m_{2}$ matrix of observed exposures. These exposures serve as an offset in the Poisson regression framework, ensuring that mortality rates are accurately scaled relative to the population at risk. Moreover, exposures can be conveniently integrated into the algorithm at a later stage, preserving the overall structure of the computational framework.
@@
The consistent aggregation structure results in a composition matrix of the form $\bm{C}=\bm{C}_{2}\otimes\bm{C}_{1}$. Each marginal composition matrix $\bm{C}_{d}$, of dimension $n_{d}\times m_{d}$ for $d=1,2$, reflects the aggregation process in its respective dimension. This tensor product formulation captures the multidimensional nature of the aggregation, ensuring that the structural relationships are preserved across all dimensions. We also assume that each latent observation only contributes to one aggregated count, i.e. the columns of the marginal composition matrices add up to one, e.g. $\bm{1}_{m_{1}}^{\prime}\bm{C}_{1}=\bm{1}^{\prime}_{n_{1}}$; this is the case in the context of mortality tables and, as we will see in the next section, a key step in the formulation of proposed approach as a Generalized Linear Array Model.
@@
Our objective is to describe the latent distribution $\bm{\gamma}$ by modeling the linear predictor $\bm{\eta}$. However, estimating a distribution with $m_{1}m_{2}$ values, which is significantly larger than the observed counts $n_{1}n_{2}$, requires imposing certain assumptions on the underlying latent structure. Given our limited understanding of this latent distribution, we propose that smoothness is a reasonable assumption for $\bm{\gamma}$, as suggested by Eilers (2007). A smooth curve inherently contains fewer details, which is appropriate given the scarcity of data. Unless there is sufficient data to support additional detailed features, smoothness allows us to extract useful information from data that might initially seem insufficient for providing any meaningful answers.
@@
where $\bm{B}_{d}$, $d=1,2$ are marginal $B$-spline basis over $\bm{x}_{d}$, of size $m_{d}\times c_{d}$, $d=1,2$. For a general reference on $B$-splines, see de Boor (2001). Currie et al. (2004) provides a clearer explanation of how to construct them in our specific context. The coefficients corresponding to each basis, denoted as $\bm{\alpha}$, must be estimated in order to derive the estimated latent distribution.
The optimal level of smoothness can be achieved either by positioning $B$-splines strategically over the covariate space or by selecting an appropriate number of $B$-splines. Alternatively, following the approach introduced by Eilers and Marx (1996) with $P$-splines, we decide to intentionally use a large number of equally-spaced $B$-splines to capture all relevant patterns in the data, then apply a penalty to remove unnecessary complexity, ensuring a parsimonious representation of the underlying structure. This translates to penalizing the coefficients $\bm{\alpha}$ in the resulting log-likelihood, a method widely employed for smoothing mortality over age and year (Currie et al., 2004; Camarda, 2019) and when dealing with grouped death counts (Rizzi et al., 2018).
@@ 2.1 Estimation procedure
The penalized log-likelihood for the Poisson PCLM in (1), where the linear predictor is defined as in (2), was presented by Eilers (2007) and holds for both one-dimensional and multidimensional cases. In both contexts, the core equation remains:
@@
where $\lambda_{1}$ and $\lambda_{2}$ control the smoothness along each respective dimension, and $\bm{D}_{d}$ denotes the difference matrix for dimension $d=1,2$ (Currie et al., 2004). In the following we use the second order differences. Different approaches can be employed to select the optimal combination of smoothing parameters $(\lambda_{1},\lambda_{2})$. However, in our context, the primary objective is to develop a more computationally efficient method for estimating a PCLM. As such, the choice of the specific selection criterion is secondary to our broader goals, and the robustness of our results will hold regardless of which criterion is ultimately chosen. We therefore subjectively select a specific combination of $(\lambda_{1},\lambda_{2})$ that demonstrates favorable performance in the illustrative applications discussed in Section 3. Any further search for the optimal smoothing parameters would only scale the computational speed arithmetically, without affecting the relative computational efficiency between the conventional and proposed approaches.
@@
$\frac{\partial\,\ell_{P}}{\partial\,\bm{\alpha}}=\bm{B}^{\prime}\bm{\Gamma}\bm{C}^{\prime}\bm{W}^{-1}\bm{y}-\bm{B}^{\prime}\exp(\bm{B}\bm{\alpha})-\bm{P}\bm{\alpha}\$ (7)
where: $\bm{W}=\verb"diag"(\bm{\mu})$ and $\bm{\Gamma}=\verb"diag"(\bm{\gamma})$. In the original work by Thompson and Baker (1981), and later extended by Eilers (2007), the solution to the system in equation (7) was approached by defining a working matrix $\breve{\bm{B}}=\bm{W}^{-1}\bm{C}\bm{\Gamma}\bm{B}$ and solving the associated GLM using standard techniques. However, as the dimensionality of the data array increases, the explicit construction of matrix $\breve{\bm{B}}$ becomes computationally prohibitive, causing a significant escalation in memory and processing requirements.
To address this issue, we propose bypassing the direct computation of $\breve{\bm{B}}$ by introducing the concept of a working latent response $\breve{\bm{y}}=\bm{\Gamma}\bm{C}^{\prime}\bm{W}^{-1}\bm{y}$, which represents the redistribution of the observed counts at the desired resolution. In fact, this would correspond to to E-step in the EM algorithm where, given current values of $\hat{\bm{\gamma}}$, $\breve{y}_{k}=\frac{\hat{\gamma}_{k}}{\sum_{l}\hat{\gamma}_{l}}y_{k}$ would approximate the unobserved latent distribution.
As we will demonstrate in the following section, defining the working latent response enables the use of a Generalized Linear Array Models (GLAM, Currie et al., 2006), which eliminates the need to compute $\breve{\bm{B}}$ or any Kronecker products involved in the estimation process. Consequently, our method is not a variant of the EM algorithm. Instead, it represents a computational strategy specifically designed to enable the estimation of a PCLM in a multidimensional framework.
@@
˘$\frac{\partial\breve{\ell}_{P}}{\partial\bm{\alpha}}=\bm{B}^{\prime}\breve{\bm{y}}-\bm{B}^{\prime}\exp(\bm{B}\bm{\alpha})-\bm{P}\bm{\alpha}\,.$ (8)
@@
RHS $=$ $-\bm{B}^{\prime}\tilde{\bm{\Gamma}}\bm{B}\tilde{\bm{\alpha}}-\bm{P}\tilde{\bm{\alpha}}-(\bm{B}^{\prime}\breve{\bm{y}}-\bm{B}^{\prime}\tilde{\bm{\gamma}}-\bm{P}\tilde{\bm{\alpha}})$
@@
where $\tilde{\bm{z}}=\tilde{\bm{\eta}}+\tilde{\bm{\Gamma}}^{-1}(\tilde{\bm{y}}-\tilde{\bm{\gamma}})$ serves a role analogous to the conventional working vector found in the GLM context, albeit evaluated at the scale of the latent response. The final system of equations presented in (9) can thus be concisely expressed as follows:
@@
This would be analogous to the M-step of the EM algorithm, where the penalized log-likelihood of the split data is maximized to obtain $\hat{\bm{\alpha}}$ (and therefore $\hat{\bm{\gamma}})$.

### v2

@@ Abstract
This paper presents a significant advancement in the estimation of the Composite Link Model within a penalized likelihood framework, specifically designed to address indirect observations of grouped count data. While the model is effective in these contexts, its application becomes computationally challenging in large, high-dimensional settings. To overcome this, we propose a reformulated iterative estimation procedure that leverages Generalized Linear Array Models, enabling the disaggregation and smooth estimation of latent distributions in multidimensional data. Through simulation studies and applications to high-dimensional mortality datasets, we demonstrate the model’s capability to capture fine-grained patterns while comparing its computational performance to the conventional algorithm. The proposed methodology offers notable improvements in computational speed, storage efficiency, and practical applicability, making it suitable for a wide range of fields in which high-dimensional data are provided in grouped formats.
Keywords: Grouped counts $\cdot\;$ Composite Link Model $\cdot\;$ Generalized Linear Array Models $\cdot\;$ Penalized splines $\cdot\;$ EM algorithm
@@ 1 Introduction
The Composite Link Model (CLM) is an advanced framework designed in situations where it is necessary to link each observation with a linear function of more than one predicted value. This is the case when the aim is to model indirect observations of counts, and this particular setting will be the focus of this paper. Initially proposed by ThompsonBakerCLM1981, the CLM extends the capabilities of the Generalized Linear Model (GLM, McCullaghNelderGLM; NelderWedderburnGLM1972). The CLM enables the incorporation of complex relationships and dependencies in count data, making it particularly useful in scenarios where direct observations are not feasible. This extension maintains the flexibility and interpretability of the GLM while addressing the unique challenges posed by indirect count data.
@@
In areas such as demography or epidemiology, this situation often arises when working with mortality tables, where death counts are aggregated over ages and/or years. Using this context as an illustration, the aim of this paper is to estimate the underlying latent mortality patterns across both ages and years, and potentially months and weeks. In the absence of prior knowledge, it is natural to exploit the inherent order in the data and impose smoothness to obtain unique estimates. This can be achieved by incorporating a roughness penalty into the associated likelihood function as described by Eilers2007. Hence, throughout this paper, we refer to it as the Penalized Composite Link Model (PCLM). The resulting scoring algorithm is a modified version of the classic iterative reweighted least-squares (IRLS) for GLMs, incorporating an additional penalty and a revised design matrix to account for the grouping structure.
While the PCLM has been successfully applied in similar contexts (RizziEtAlPCLM1D2015; RizziEtAlPCLM2D2018), including within Bayesian estimation frameworks (LambertEilersBayesGrouped2009; LambertBayesBivariateCSDA2011), the problem size in two- or higher-dimensional settings often becomes quite large, rendering the computational cost of maximizing the penalized likelihood function infeasible. Recently, deep learning approaches have been proposed for ungrouping data across multiple populations, a setting that can be viewed as moderately high-dimensional (NigriEtAlDisaggregating2024). These methods, however, are tailored to mortality data and lie outside the PCLM framework: they do not model the likelihood of the observed aggregated counts and therefore do not allow for likelihood-based inference. In contrast, this paper focuses on the PCLM itself and proposes an alternative estimation procedure that enables its efficient estimation in genuinely high-dimensional settings, based on a reformulation of IRLS that resembles the EM-algorithm (DempsterEtAl_EM1977). We alternate between computation of the latent distribution of counts from the current parameter estimates and estimating the parameters from the actual and estimated observations. Although it is well established that the EM algorithm can suffer from slow convergence (NgEtAl_EMchapter2012), the computational and storage advantages of our approach stem from expressing the whole algorithm as a Generalized Linear Array Model (GLAM, CurrieDurbanEilersGLAM2006). This formulation avoids the need to explicitly construct large and memory-intensive model matrices and efficiently handles the large systems of equations that arise. As a result, our method achieves substantial gains in both memory usage and computational speed. Furthermore, unlike the EM algorithm, which assumes a Poisson distribution for latent variables rather than for the observed aggregated counts, our approach preserves the original probabilistic assumptions of the model. It does not alter the underlying data-generating process, but instead constitutes a re-engineering of the estimation procedure to accommodate high-dimensional settings efficiently.
This methodology could be extremely useful in common situations where the grouping structure in one dimension is constant across the other dimension (this is a common case in mortality tables where we observed aggregated death counts in coarse age groups and the aggregation structure is similar for all years and other dimensions). Although exact uniform grouping may not always be observed in real-world datasets, the assumption of consistent aggregation is often approximately satisfied in many practical contexts, particularly in demography and spatio-temporal epidemiology. For instance, in spatio-temporal health surveillance, data are commonly collected at regular temporal intervals (e.g., weekly counts), even when spatial granularity is heterogeneous or sparse. In such cases, the temporal dimension typically imposes a structured and regular grouping pattern (Leeetal2022). Furthermore, aggregation structures in space-time settings often exhibit a regular and partially separable structure across dimensions. This characteristic enables them to be effectively represented using combinations of their marginal temporal and spatial components that are well aligned with the assumptions of our proposed modeling framework. As a result, the methodology retains broad applicability, particularly in settings where at least one dimension exhibits consistent and well-defined grouping patterns.
In summary, the main contributions of this work can be summarized in five points. First, a new computational formulation for Penalized Composite Link Models is introduced, which avoids the explicit construction of high-dimensional Kronecker-product design and composition matrices, a central feature of existing implementations. Second, this formulation enables the estimation of PCLMs in two or more dimensions at fine resolutions that are computationally infeasible or unattainable with currently available methods. Third, the original probabilistic assumptions of the Composite Link Model are fully preserved, avoiding the additional modeling assumptions and likelihood modifications inherent to EM-based approaches. Fourth, it allows full variance and uncertainty estimation in multidimensional settings, where existing algorithms frequently fail due to prohibitive memory requirements. Finally, the approach achieves substantial gains in computational speed and memory efficiency without compromising estimation accuracy.
@@
The structure of this paper is as follows: following this Introduction, Section 2 provides an overview of the Penalized Composite Link Model, focusing on the new estimation methodology within the Generalized Linear Array Models framework and on methods for computing variance and standard errors. Section 3 presents two simulation studies designed to evaluate the accuracy of the PCLM in recovering the latent distribution and to compare the computational performance of the proposed method against the original formulation. Section 4 illustrates the practical utility of the proposed approach through the analysis of two mortality datasets in two-dimensional and three-dimensional settings. Finally, Section 5 concludes by summarizing the main findings.
@@ 2 The Penalized Composite Link Model over multidimensional arrays
Suppose that we observe an array of counts $\bm{Y}=(y_{i,j})$ (of size $n_{1}\times n_{2}$), where $n_{1}$ and $n_{2}$ correspond to the number of coarse-level observations along two covariates: $\bm{x}_{1}=(x_{11},\ldots,x_{1n_{1}})$ and $\bm{x}_{2}=(x_{21},\ldots,x_{2n_{2}})$ representing the grid over which the observations are aggregated. We assume that $Y_{i,j}$, the random variable corresponding to the observed counts $y_{i,j}$, follows the Poisson distribution with mean $\mu_{i,j}$, i.e.  $Y_{i,j}\sim\mathcal{P}(\mu_{i,j})$. Let $\bm{y}=\verb"vec"(\bm{Y})$ denote the vectorized form of the observed counts (of length $n_{1}n_{2}$), and $\bm{\mu}$ represent the corresponding vector of expected values. In our setting, observed counts are the result of the contribution of several latent observations which we aim to estimate over a finer resolution of the covariates $\bm{x}_{1}$ and $\bm{x}_{2}$. In other words, while we observe $\bm{y}$, we aim to estimate the vector of latent observations $\bm{\gamma}=(\gamma_{11},\gamma_{12},\ldots,\gamma_{m_{1}m_{2}})$, where $m_{1}>n_{1}$ and $m_{2}>n_{2}$ denote the number of fine-scale levels along each covariate. The relation between the latent vector $\bm{\gamma}$ and $\bm{\mu}$ is specified by a known composition matrix $\bm{C}$ (of size $n_{1}n_{2}\times m_{1}m_{2}$), which describes how the elements of the latent vector $\bm{\gamma}$ are combined to yield $\bm{\mu}$. The resulting Composite Link Model for Poisson aggregated counts can thus be written as follows:
@@
For our illustrative example, we will utilize demographic data on mortality. Here, the counts represent observed deaths, with $\bm{x}_{1}$ corresponding to ages and $\bm{x}_{2}$ to calendar years. In many cases, especially when dealing with smaller geographic areas or specific causes of death, the available data are often aggregated over broader age groups or multiple years. Our objective is to estimate mortality at a more granular level, specifically for individual ages and single calendar years, to provide finer insights into mortality patterns. In general, it is common for the aggregation pattern in each dimension of the array to be consistent across all dimensions. For instance, if death counts are aggregated into 5-year age intervals, this grouping pattern is typically applied systematically across all (grouped) years of death. Furthermore, when analyzing death counts, it is essential to recognize that the expected values in the Poisson distribution are the product of the underlying force of mortality and the population at risk, commonly referred to as exposures: $\bm{y}\sim\mathcal{P}(\bm{e}\odot\bm{\mu})$. Exposures can be simply multiplied by the underlying latent force of mortality over a finer resolution since they are typically provided for single years of age and time, i.e. $\bm{e}$ represents the vectorized form of the $m_{1}\times m_{2}$ matrix of observed exposures. These exposures serve as an offset in the Poisson regression framework, ensuring that mortality rates are accurately scaled relative to the population at risk. Moreover, exposures can be conveniently integrated into the algorithm at a later stage, preserving the overall structure of the computational framework.
@@
The consistent aggregation structure results in a composition matrix of the form $\bm{C}=\bm{C}_{2}\otimes\bm{C}_{1}$. Each marginal composition matrix $\bm{C}_{d}$, of dimension $n_{d}\times m_{d}$ for $d=1,2$, reflects the aggregation process in its respective dimension. This tensor product formulation captures the multidimensional nature of the aggregation, ensuring that the structural relationships are preserved across all dimensions. We also assume that each latent observation only contributes to one aggregated count, i.e. the columns of the marginal composition matrices add up to one, e.g. $\bm{1}_{m_{1}}^{\prime}\bm{C}_{1}=\bm{1}^{\prime}_{n_{1}}$; this is the case in the context of mortality tables and, as we will see in the next section, a key step in the formulation of the proposed approach as a Generalized Linear Array Model.
@@
Our objective is to describe the latent distribution $\bm{\gamma}$ by modeling the linear predictor $\bm{\eta}$. However, estimating a distribution with $m_{1}m_{2}$ values, which is significantly larger than the observed counts $n_{1}n_{2}$, requires imposing certain assumptions on the underlying latent structure. Given our limited understanding of this latent distribution, we propose that smoothness is a reasonable assumption for $\bm{\gamma}$, as suggested by Eilers2007. A smooth curve inherently contains fewer details, which is appropriate given the scarcity of data. Unless there are sufficient data to support additional detailed features, smoothness allows us to extract useful information from data that might initially seem insufficient for providing any meaningful answers.
@@
where $\bm{B}_{d}$, $d=1,2$ are marginal $B$-spline bases over $\bm{x}_{d}$, of size $m_{d}\times c_{d}$, $d=1,2$. For a general reference on $B$-splines, see Boor2001. CurrieDurbanEilersSmoothForecast2004 provides a clearer explanation of how to construct them in our specific context. The coefficients corresponding to each basis, denoted as $\bm{\alpha}$, must be estimated in order to derive the estimated latent distribution.
The optimal level of smoothness can be achieved either by positioning $B$-splines strategically over the covariate space or by selecting an appropriate number of $B$-splines. Alternatively, following the approach introduced by EilersMarx1996 with $P$-splines, we decide to intentionally use a large number of equally-spaced $B$-splines to capture all relevant patterns in the data, then apply a penalty to remove unnecessary complexity, ensuring a parsimonious representation of the underlying structure. This translates to penalizing the coefficients $\bm{\alpha}$ in the resulting log-likelihood, a method widely employed for smoothing mortality over age and year (CurrieDurbanEilersSmoothForecast2004; CamardaCPsplines2019) and when dealing with grouped death counts (RizziEtAlPCLM2D2018).
@@ 2.1 Estimation procedure
The penalized log-likelihood for the Poisson PCLM in (1), where the linear predictor is defined as in (2), was presented by Eilers2007 and holds for both one-dimensional and multidimensional cases. In both contexts, the core equation remains:
@@
where $\lambda_{1}$ and $\lambda_{2}$ control the smoothness along each respective dimension, and $\bm{D}_{d}$ denotes the difference matrix for dimension $d=1,2$ (CurrieDurbanEilersSmoothForecast2004). In the following, we use the second order differences. Different approaches can be employed to select the optimal combination of smoothing parameters $(\lambda_{1},\lambda_{2})$. However, in our context, the primary objective is to develop a more computationally efficient method for estimating a PCLM. As such, the choice of the specific selection criterion is secondary to our broader goals, and the robustness of our results will hold regardless of which criterion is ultimately chosen. We therefore subjectively select a specific combination of $(\lambda_{1},\lambda_{2})$ that demonstrates favorable performance in the illustrative applications discussed in Section 4. Any further search for the optimal smoothing parameters would only scale the computational speed arithmetically, without affecting the relative computational efficiency between the conventional and proposed approaches.
@@
$\frac{\partial\,\ell_{P}}{\partial\,\bm{\alpha}}=\bm{B}^{\prime}\bm{\Gamma}\bm{C}^{\prime}\bm{W}^{-1}\bm{y}-\bm{B}^{\prime}\exp(\bm{B}\bm{\alpha})-\bm{P}\bm{\alpha}\\$ (7)
where: $\bm{W}=\verb"diag"(\bm{\mu})$ and $\bm{\Gamma}=\verb"diag"(\bm{\gamma})$. In the original work by ThompsonBakerCLM1981, and later extended by Eilers2007, the solution to the system in equation (7) was approached by defining a working matrix $\breve{\bm{B}}=\bm{W}^{-1}\bm{C}\bm{\Gamma}\bm{B}$ and solving the associated GLM using standard techniques. However, as the dimensionality of the data array increases, the explicit construction of matrix $\breve{\bm{B}}$ becomes computationally prohibitive, causing a significant escalation in memory and processing requirements.
To address this issue, we propose bypassing the direct computation of $\breve{\bm{B}}$ by introducing the concept of a working latent response $\breve{\bm{y}}=\bm{\Gamma}\bm{C}^{\prime}\bm{W}^{-1}\bm{y}$, which represents the redistribution of the observed counts at the desired resolution. In fact, this would correspond to the E-step in the EM algorithm where, given current values of $\hat{\bm{\gamma}}$, $\breve{y}_{k}=\frac{\hat{\gamma}_{k}}{\sum_{l}\hat{\gamma}_{l}}y_{k}$ would approximate the unobserved latent distribution.
As we will demonstrate in the following section, defining the working latent response enables the use of a Generalized Linear Array Model (GLAM, CurrieDurbanEilersGLAM2006), which eliminates the need to compute $\breve{\bm{B}}$ or any Kronecker products involved in the estimation process. Consequently, our method is not a variant of the EM algorithm. Instead, it represents a computational strategy specifically designed to enable the estimation of a PCLM in a multidimensional framework.
@@
$\frac{\partial\breve{\ell}_{P}}{\partial\bm{\alpha}}=\bm{B}^{\prime}\breve{\bm{y}}-\bm{B}^{\prime}\exp(\bm{B}\bm{\alpha})-\bm{P}\bm{\alpha}\,.$ (8)
@@
[…]
@@
[…]
@@
[…]